OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

EQUATION OF GIVEN LINE IS y=-x-3 its slope is -1 thus aline perpendicular to it has a slope = -1/-1 = 1 now the point is (2,3) using point slope for we get y-y1=m(x-x1) y-3= 1(x-2) =>y=x-2+3 =>y=x+1------------>required equation

OpenStudy (anonymous):

thanks!